# Fault Diagnosis and Prognosis Capabilities for Wind Turbine Hydraulic Pitch Systems: An Overview

### v1

@@ Abstract
Wind energy is the leading non-hydro renewable technology. Increasing reliability is a key factor in reducing the downtime of high-power wind turbines installed in remote off-shore places, where maintenance is costly and less reactive. Defects in the pitch system are responsible for up to $20\%$ of a wind turbine downtime. Thus, prognosis and condition monitoring of such defects are essential for avoiding downtime. This paper presents a generic assessment of diagnosis capabilities in hydraulic pitch systems characterizing high-power wind turbines. A mathematical model of the non-linear system dynamics is presented along with a description of the most frequent faults that can affect it. Structural analysis is employed to assess the diagnosis and prognosis capabilities of defects in the pitch system. The structural properties are furthermore explored to investigate possible sensor reduction without compromising fault diagnosis capabilities.
@@ 1 Introduction
Wind turbines production increased by a record of 273 TWh ($17\%$) in 2021222Data of 2022 is less meaningful because of the energy crisis, more details at (ieaWorldEnergyOutlook2022, )., the highest among all renewable power technologies. The same year the total production reads 1870 TWh, almost as much as all the other non-hydro renewable technologies combined. The Net Zero Emissions target by 2050 is reflected in the tendency to increase the turbine’s size and power and to move towards off-shore systems(ieaWindElectricityAnalysis2022, ). Minimizing the downtime for such systems is a priority. In addition to the maintenance cost, one needs to consider the missing revenue associated with the energy not being produced during the repair. The offshore installation makes maintenance more challenging and, thus, longer in time. Studies like (gayoFinalReportSummary2011, ; carrollFailureRateRepair2016, ; linigerReliableFluidPower2015, ) have shown that the main source of failure in wind turbines is associated with the pitch regulation system. The authors in (chengInfluenceMechanicalFaults2023, ), (etemaddarResponseAnalysisComparison2016a, ) and (leeFailureAnalysisHydraulic2020a, ) have shown how failures in such systems may have severe consequences. For high-power models, the pitch regulation system is electro-hydraulic. Compared to purely electrical pitch systems, electro-hydraulic systems offer higher power density, more robustness, and smoother performances (walgernReliabilityElectricalHydraulic2023, ; palavicinoFaultsDiagnosticsElectrified2018, ; luReviewRecentAdvances2009, ). The application of fault detection and accommodation schemes can be leveraged to reduce the required maintenance on the hydraulic subsystem such that the up-time of the wind turbine significantly increases.
@@
A significant corpus of literature has been focusing on the design and evaluation of methods for the detection and mitigation of individual faults that can occur in wind turbines. A comprehensive review of the research methods was presented in (badihiComprehensiveReviewSignalBased2022, ; fekihFaultDiagnosisFault2022, ). Actuator faults have also been addressed in the context of fault-tolerant control in (songFaultTolerantControlFloating2023, ; palanimuthuReliabilityImprovementLargescale2023, ; liuActuatorFaultTolerant2023, ; mousaviFaulttolerantOptimalPitch2022, ). The authors in (asmussenFaultDetectionDiagnosis2021, ) focused on the hydraulic pitch system and reviewed most of the results for both fault diagnosis and fault-tolerant control. Other investigations related to fault diagnosis and mitigation in the pitch system can be found in (elorzaSensorDataProcessing2022a, ; venkaiahPitchControlElectrohydraulic2023, ; lanReviewFaultDiagnosis2023a, ) and the references therein. The hydraulic circuit driving the pitch system in wind turbines is similar to the ones used in several other applications in hydraulics, such as mechanical presses, servo positioning, etc. Several studies applied similar tools to detect and mitigate mainly leakage faults. The most recent and relevant results for detection/estimation and mitigation can be found in (goharriziWaveletBasedApproachExternal2011, ; vasquezActiveFaultDiagnosis2019, ; aslAdaptiveSquarerootUnscented2019, ; bahramiAdaptiveSupertwistingObserver2018a, ; shanbhagFailureMonitoringPredictive2021, ; choFaultDetectionDiagnosis2021, ; zhaoFaultDiagnosisHydraulic2020, ; liuFaultDiagnosisHydraulic2014, ; goharriziInternalLeakageDetection2012, ; xuModelBasedFaultDetection2015, ; sharifiMulticlassFaultDetection2016, ; sunMultiFaultDiagnosisApproach2020, ; zhangSARPerformancebasedFault2022, ; djordjevicSensorFaultEstimation2022, ; daiSignalBasedIntelligentHydraulic2019, ; shenFaultDiagnosisAircraft2022, ) and (maddahiPracticalApproachDesigning2020, ; daoActiveFaultTolerant2021, ; phanOptimizedBasedFaultTolerantControl2022a, ; phanActuatorFailureCompensationbased2023, ; phanFaulttolerantControlElectrohydraulic2023, ; zhangMultiModelBasedAdaptive2022, ), respectively. The references include results for model-based, statistical-based, and learning-based methods.
The majority of the previously reported work pertained to the design of fault-tolerant solutions in relation to specific faults without considering the overall fault diagnosis capabilities of the system. To the best of the authors’ knowledge, such a holistic investigation has not been carried out. This study focuses on providing a means to systematically assess the potential fault tolerance capabilities of hydraulic pitch systems. Specifically, the contributions of this paper are:
@@
    Derivation of a benchmark model including explicit representations of the most relevant faults.
@@
    Structural analysis of the benchmark model for assessing the fault diagnosis capabilities of the system.
@@
    Investigation of sensor topology in connection to cost reduction while retaining fault diagnosis features.
The remainder of the paper is structured as follows: Section 2 defines a benchmark model that is used for investigating condition monitoring and fault diagnosis in hydraulic pitch systems. The fault detection and isolation capabilities of the system are analyzed in Section 3 by carrying out a structural analysis on the benchmark model. Section 4 investigates topology alterations that could improve fault diagnosis in the system. The results are discussed in Section 5, and conclusions are drawn in Section 6.
@@ 2 Pitch System Topology and Modeling
The reference hydraulic pitch system for the article is presented in Figure 1 and represents a simplified schematics/topology of a pitch system under normal operation, but excluding the safety features. A pump and an accumulator bank (represented as one accumulator) are used to supply three equivalent circuits composed of a proportional valve (PV) and a hydraulic cylinder. More information about how such a system works is provided in (linigerDesignReliableFluid2018, ).
Figure 1: Hydraulic pitch system components (pedersenInvestigationLoadReduction2016, ).
The benchmark model comprises the nominal dynamics detailed in (pedersenInvestigationLoadReduction2016, ) and all major faults that could affect turbine availability and production. Although modeling for fault-tolerance may require a multiplicative representation of some defects (blankeDiagnosisFaultTolerantControl2016, ), all faults in this study are expressed as additive, time-varying terms (in blue color) since the effect of any generic fault can be described in this way. A thorough description and models for diagnosis are presented in Section 2.5. The sections  2.1 to  2.3 present the mathematical model for the different elements comprised in the pitch system.
@@ 2.1 Hydraulic Cylinder
The main component allowing the pitching of the blades is the hydraulic cylinder. Its dynamics is described by the force balance: {linenomath}
$M_{eq}\dot{v_{c}}=A_{p}p_{p}-A_{r}p_{r}-B_{v}v_{c}-F_{c}\,tanh(v_{c}/\gamma)\,{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}+f_{Fr,c}}-F_{ext}$ (1)
@@
$\beta_{e}=\frac{1}{\frac{1}{\beta_{oil}}+\varepsilon_{a}(p)\left(\frac{1}{c_{ad}p}-\frac{1}{\beta_{oil}}\right)}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}+f_{B,e}}$ (4)
$\varepsilon_{a}(p)=\frac{1}{\frac{1-\varepsilon_{a,0}}{\varepsilon_{a,0}}\left(\frac{p_{atm}}{p}\right)^{\frac{-1}{c_{ad}}}+1}.$ (5)
Its value differs for different cylinder chambers and for the supply circuit. The fault $f_{B,e}$ is related to the oil degradation, regardless of the part of the plant where it first appears. For this reason, a unique value for $\beta_{e}$ is used to describe the general effective bulk modulus, which depends on the supply pressure $p_{s}$, introduced further down. An example of the effective bulk modulus dependency on air content, pressure, and temperature is described in (knezevicAnalysisChangesBulk2011, ). Of the three, the air content is by far the most significant, but as the system normally will be operating with pressures well above 25 bar, changes in the effective bulk modulus will be very limited, even for air content up to 10% volumetric air content (measured at atmospheric pressure). Therefore, the model does not consider temperature as a state variable. However, it is probably unlikely to detect a fault in the bulk modulus in practice because of all the uncertainties associated with the available models. Nevertheless, it is reported in the complete model to highlight its importance.
@@
$Q_{le,p}=C_{le,p}(p_{p}-p_{atm}){\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}+f_{Q_{le,p}}}$ (6)
$Q_{le,r}=C_{le,r}(p_{r}-p_{atm}){\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}+f_{Q_{le,r}}}$ (7)
$Q_{li}=C_{li}(p_{p}-p_{r}){\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}+f_{Q_{li}}}.$ (8)
@@ 2.2 Valves
$\left.-\sqrt{|p_{p}-p_{t}|}sign(p_{p}-p_{t})H(-x_{v})\right]{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}+f_{Q_{p}}}$ (10)
$Q_{rv}=-K_{v}(x_{v})\phi_{v}\sqrt{|p_{s}-p_{r}|}sign(p_{s}-p_{r})H(-x_{v}){\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}+f_{Q_{rv}}},$ (11)
where $H(\cdot)$ is the Heaviside step function for distinguishing the different results for the different operating regions of the valve, i.e.,
@@
$\dot{v_{v}}+2\xi\omega_{0}\,{v_{v}}+\omega_{0}^{2}\,x_{v}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}+f_{Fr,v}}=\omega_{0}^{2}\,x_{v,ref}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}+f_{wv,v}},\quad x_{v,ref}=k_{u}u.$ (13)
@@
$Q_{cv}=K_{cv}(p_{r}-p_{s}-p_{cv,c})H(p_{r}+p_{cv,c}-p_{s}){\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}+f_{Q_{cv}}}.$ (15)
@@ 2.3 Supply Circuit
$Q_{acc}=Q_{s}-Q_{rel}+\sum_{i=1}^{3}Q_{cv,i}-\sum_{i=1}^{3}Q_{v,i}\,{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}+f_{Q_{ru}}}.$ (17)
@@
$Q_{rel}=K_{rel}(p_{s}-p_{cr,r})H(p_{s}-p_{cr,r}){\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}+f_{Q_{rel}}}.$ (19)
@@
$V_{gas}=V_{acc}\left[\left(\frac{p_{gas,0}}{p_{s}}\right)^{1/k}H(p_{s}-p_{gas,0})\right]{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}+f_{acc}}.$ (21)
@@ 2.5 Fault Description
In the previous subsections, faults have been inserted as additive terms in the equations describing the nominal plant operation. For setting up a structural analysis, it is important to specify where the fault enters, as that is the only relevant information for evaluating whether it is possible to detect it or isolate it from other faults. In reality, since different faults have different sources, they also have different properties as signals. This information is only relevant in the design phase of the diagnostic tools. Table 1 presents a basic description of all the faults according to (papageorgiouOnlineFrictionParameter2020, ; asmussenFaultDetectionDiagnosis2021, ; choNumericalModelingHydraulic2020, ; linigerReliabilityBasedDesign2017, ; shanbhagFailureMonitoringPredictive2021, ; khanFaultDetectionElectrohydraulic2002, ).
Table 1: Failure Modes. Please notice that the rating of the failures is based on that a certain threshold is exceeded for each type of failure, where the failure becomes significant.
Fault Name Fault Model Description Severity $f_{Fr,c}$ $\Delta_{B_{c}}v_{c}+\Delta_{F_{c}}\,tanh(v_{c}/\gamma)$ Increased friction in the cylinder because of sludge formation. Not critical under a certain threshold, but it can cause the cylinder to get stuck. In the initial phase it appears as a variation of friction coefficients. High $f_{B,e}$ $\Delta_{B,e}$ Fault related to oil (i.e. gas) contamination, difficult to model differently than a generic variation with respect to the nominal value. Medium $f_{Q_{le,p}}\,f_{Q_{le,r}}$ $\Delta_{C_{le,i}}(p_{i}-p_{atm})$ Oil leakage between a cylinder chamber and the outside. It can be slow or abrupt, according to wear. Even with small values the turbine will be shut down. From a control point of view, leakage do however act to increase the damping in the system. High $f_{Q_{li}}$ $\Delta_{C_{li}}(p_{p}-p_{r})$ Oil leakage between the two chambers in the cylinder. Same characteristics as the external one, it just increases damping in the pitching. Low $f_{Q_{p}}\,f_{Q_{rv}}$ $\Delta_{K_{v}}[\cdot]$ Variation in the proportionality coefficient due to mechanical wear of the valve, or by incorrect valve command. As long as the valve is able to operate the severity is low. Low $f_{Fr,v}$ $2\Delta_{\xi}\Delta_{\omega_{0}}\,{v_{v}}+\Delta_{\omega_{0}}^{2}\,(x_{v}-x_{v,ref})$ Emergence of significant friction in the valve because of sludge formation or wear. The effect is seen in the closed-loop dynamics parameters, and may lead to the valve getting stuck with time. High $f_{wv,v}$ $\omega_{0}^{2}\Delta_{k_{u}}u$ Fault in the coil, leading to a wrong voltage setting. It is modeled as a wrong position setting since the impact is equivalent. High $f_{Q_{cv}}$ $\Delta_{Q_{cv}}$ It’s usually a discrete-time nature fault leading to the valve getting stuck, more than leakages or a variation in the proportional flow coefficient. High $f_{Q_{ru}}$ $\Delta_{Q_{ru}}$ Leakage flow in the rotary unit, i.e., the connection between the pressure source and the three actuation circuits. Low $f_{Q_{rel}}$ $\Delta_{Q_{rel}}$ It can be a variation in the proportional coefficient, leakage, or a variation in the pressure at which it is triggered, although these type of faults are rare. Low $f_{acc}$ $-V_{gas,l}$ Gas leakage in the accumulator, with consequent reduction of the volume of gas, for a given pressure. Generally slowly varying. Medium $f_{y,p_{s}}$$f_{y,p_{p}}$$f_{y,p_{r}}$ $\Delta_{y,p_{i}}$ Wrong measurement in the related pressure sensor. Low $f_{y,x_{c}}$$f_{y,x_{v}}$ $\Delta_{y,x_{i}}$ Wrong measurement in the cylinder/valve position sensor. Position measurements are the only signals used in the control loop. Hence, their importance. High
The fault description column provides information on how different faults are related to each other. The severity of a given fault depends on how it propagates in the system. Figure 2 illustrates such propagation properties of the considered faults.
@@
The mathematical model and Table 1 do not include any description related to the pump failure, as there is no dynamical relation associated with it. A fault in the pump will, from an operational point of view, typically result in reduced pump flow and ripples, leading to slower pressure build-up in the system. One means of detecting such a fault is to monitor the pump pressure by installing a pressure sensor in the vicinity of the pump.

### v2

@@ Abstract
Wind energy is the leading non-hydro renewable technology. Increasing reliability is a key factor in reducing the downtime of high-power wind turbines installed in remote off-shore places, where maintenance is costly and less reactive. Defects in the pitch system are responsible for up to $20\%$ of a wind turbine downtime. Thus, monitoring such defects is essential for avoiding it. This paper presents a generic assessment of the diagnosis capabilities in hydraulic pitch systems, which are used in high-power wind turbines. A mathematical model of the non-linear system dynamics is presented along with a description of the most frequent faults that occur. Structural analysis is used to assess which defects can be detected in the pitch system. The structural properties are furthermore explored to investigate the possibility of reducing the amount of sensors without compromising the fault diagnosis capabilities. Robustness to model uncertainty is finally addressed and generic principles for estimating the detectable magnitude of wear and tear are presented.
@@ 1 Introduction
Wind turbines production increased by a record of 273 TWh ($17\%$) in 2021222Data of 2022 is less meaningful because of the energy crisis, more details at [29]., the highest among all renewable power technologies. The same year the total production reads 1870 TWh, almost as much as all the other non-hydro renewable technologies combined. The Net Zero Emissions target by 2050 is reflected in the tendency to increase the turbine’s size and power and to move towards off-shore systems [28]. Minimizing the downtime for such systems is a priority. In addition to the maintenance cost, one needs to consider the missing revenue associated with the energy not being produced during the repair. The offshore installation makes maintenance more challenging and, thus, longer in time. Studies like [24, 8, 37] have shown that the main source of failure in wind turbines is associated with the pitch regulation system. The authors in [10], [20] and [35] have shown how failures in such systems may have severe consequences. For high-power models, the pitch regulation system is electro-hydraulic. Compared to purely electrical pitch systems, electro-hydraulic systems offer higher power density, more robustness, and smoother performances [62, 47, 42]. The application of fault detection and accommodation schemes can be leveraged to reduce the required maintenance on the hydraulic subsystem such that the up-time of the wind turbine significantly increases.
@@
A significant corpus of literature has been focusing on the design and evaluation of methods for the detection and mitigation of individual faults that can occur in wind turbines. A comprehensive review of the research methods is presented in [3, 21, 22, 63, 65]. Actuator faults have also been addressed in the context of fault-tolerant control in [57, 46, 41, 44]. The authors in [2] focused on pitch system defects. It reviewed results from pitch system fault detection and also mentioned possibilities with fault-tolerant control. Other investigations related to fault diagnosis and mitigation for the pitch system included [19, 61, 33] and the references therein. The hydraulic circuit driving the pitch system in wind turbines is similar to the ones used in several other applications for heavy machinery, such as mechanical presses and servo control of ship’s rudders and variable pitch propellers. Several studies applied tools to detect and mitigate mainly leakage faults. The most recent and relevant results for detection/estimation and mitigation appear in [26, 60, 1, 4, 54, 12, 68, 40, 27, 64, 55, 58, 66, 16, 13, 56, 25, 59] and [43, 15, 50, 52, 51, 67]. The above references consider results for model-based, statistical-based, and learning-based methods, with focus on specific selected defects.
Employing graph-based methods of topology, this paper provides a holistic investigation that systematically assess capability detecting and isolate defects in a hydraulic pitch system, and also addressing what is the coverage of diagnosis with the typical sensor configuration in a high-power wind turbine. The contributions of the paper are:
@@
    Include faults essential for resilience and maintenance in a generic benchmark model.
@@
    Employ structural analysis of the model to assess generic capabilities for diagnosis.
@@
    Investigate sensor topology aiming at cost reduction while retaining coverage of diagnosis.
  • 4.
    Assess the robustness of fault detection to model uncertainty.
The remainder of the paper is structured as follows: Section 2 defines a benchmark model that is used to investigatecondition monitoring and fault diagnosis in hydraulic pitch systems. The fault detection and isolation capabilities of the system are analyzed in Section 3 by carrying out a structural analysis on the benchmark model. Section 4 investigates topology alterations that could improve fault diagnosis in the system. Robustness of fault detection with respect to model uncertainty is treated in Section 5. The results are discussed in Section 6, and conclusions are drawn in Section 7.
@@ 2 Pitch System Topology and Modeling
The reference hydraulic pitch system for the article is presented in Figure 1 and represents a simplified schematics/topology of a pitch system under normal operation, but excluding the safety features. A pump and an accumulator bank (represented as one accumulator) are used to supply three equivalent circuits composed of a proportional valve (PV) and a hydraulic cylinder. More information about how such a system works is provided in [36].
Figure 1: Hydraulic pitch system components [49].
The benchmark model comprises the nominal dynamics detailed in [49] and all major faults that could affect turbine availability and production. Although modeling for fault-tolerance may require a multiplicative representation of some defects [5], all faults in this study are expressed as additive, time-varying terms (in blue color) since the effect of any generic fault can be described in this way when considering diagnosis. A thorough description and models for diagnosis are presented in Section 2.5, and sections  2.1 to  2.3 present the detailed mathematical models for the elements of the pitch system.
@@ 2.1 Hydraulic Cylinder
The main component allowing the pitching of the blades is the hydraulic cylinder. Its dynamics is described by the force balance:
$M_{eq}\dot{v_{c}}=A_{p}p_{p}-A_{r}p_{r}-B_{v}v_{c}-F_{c}\,tanh(v_{c}/\gamma)\,{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}+f_{Fr,c}}-F_{ext}$ (1)
@@
$\beta_{e}=\frac{1}{\frac{1}{\beta_{oil}}+\varepsilon_{a}(p)\left(\frac{1}{c_{ad}p}-\frac{1}{\beta_{oil}}\right)}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}+f_{B,e}}$ (4)
$\varepsilon_{a}(p)=\frac{1}{\frac{1-\varepsilon_{a,0}}{\varepsilon_{a,0}}\left(\frac{p_{atm}}{p}\right)^{\frac{-1}{c_{ad}}}+1},$ (5)
where the fault $f_{B,e}$ is related to oil degradation.
An example of the effective bulk modulus dependency on air content, pressure, and temperature is described in [31]. Of the three, the air content is by far the most significant, but as the system normally will be operating with pressures well above 25 bar, changes in the effective bulk modulus will be very limited, even for air content up to 10% volumetric air content (measured at atmospheric pressure). Therefore, the model does not consider temperature as a state variable.
The high degree of uncertainty in the bulk modulus model described in Equations 4 and 5, and the fact that the pressures in the cylinders’ chambers as well as in the supply circuit vary within similar values, has motivated the adoption of a single uncertain parameter $\beta_{e}$ to describe the effective bulk modulus in the entire circuit. This simplification is consistent with the fact that the fault $f_{B,e}$ is related to the oil degradation, regardless of the part of the plant where it appears. Thus, the simplified effective bulk modulus description is givne by the following equation:
$\beta_{e}=\beta_{e,0}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}+f_{B,e}}$ (6)
@@
$Q_{le,p}=C_{le,p}(p_{p}-p_{atm}){\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}+f_{Q_{le,p}}}$ (7)
$Q_{le,r}=C_{le,r}(p_{r}-p_{atm}){\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}+f_{Q_{le,r}}}$ (8)
$Q_{li}=C_{li}(p_{p}-p_{r}){\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}+f_{Q_{li}}}.$ (9)
@@ 2.2 Valves
$\left.-\sqrt{|p_{p}-p_{t}|}sign(p_{p}-p_{t})H(-x_{v})\right]{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}+f_{Q_{p}}}$ (11)
$Q_{rv}=-K_{v}(x_{v})\phi_{v}\sqrt{|p_{s}-p_{r}|}sign(p_{s}-p_{r})H(-x_{v}){\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}+f_{Q_{rv}}},$ (12)
where $K(x_{v})$ is the tabular function with which the manufacturer describes the input-to-flow relation of the controllable valve.The Heaviside step function $H(\cdot)$ distinguishes the different operating regions of the valve, i.e.,
@@
$\dot{v_{v}}+2\xi\omega_{0}\,{v_{v}}+\omega_{0}^{2}\,x_{v}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}+f_{Fr,v}}=\omega_{0}^{2}\,x_{v,ref}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}+f_{wv,v}},\quad x_{v,ref}=k_{u}u.$ (14)
@@
$Q_{cv}=K_{cv}(p_{r}-p_{s}-p_{cv,c})H(p_{r}-p_{s}-p_{cv,c}){\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}+f_{Q_{cv}}}.$ (16)
@@ 2.3 Supply Circuit
$Q_{acc}=Q_{s}-Q_{rel}+\sum_{i=1}^{3}Q_{cv,i}-\sum_{i=1}^{3}Q_{v,i}\,{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}+f_{Q_{ru}}}.$ (18)
@@
$Q_{rel}=K_{rel}(p_{s}-p_{cr,r})H(p_{s}-p_{cr,r}){\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}+f_{Q_{rel}}}.$ (20)
@@
$V_{gas}=V_{acc}\left[\left(\frac{p_{gas,0}}{p_{s}}\right)^{1/k}H(p_{s}-p_{gas,0})\right]{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}+f_{acc}}.$ (22)
@@ 2.5 Fault Description
In the previous subsections, faults have been inserted as additive terms in the equations describing the nominal plant operation. For setting up a structural analysis, it is important to specify where the fault enters, as that is the only relevant information for evaluating whether it is possible to detect it or isolate it from other faults. In reality, since different faults have different sources, they also have different properties as signals. This information is only relevant in the design phase of the diagnostic tools. Table 1 presents a basic description of all the faults according to [48, 2, 11, 39, 54, 30].
The description column provides information on how different faults are related. The severity of a given fault depends on how it propagates in the system. Figure 2 illustrates such propagation properties of the faults listed in the table.
The mathematical model and Table 1 do not include description related to pump failure, as there is no dynamical relation associated with it. A fault in the pump will, from an operational point of view, typically result in reduced pump flow and ripples, leading to slower pressure build-up in the system. Detection of such a fault is obtained by pump pressure monitoring using a pressure sensor in the vicinity of the pump.
Table 1: Failure Modes. The rating of failures when a component specific threshold is exceeded.
[…]